OpenStudy (anonymous):

x^2 +18x +? Complete the square

OpenStudy (anonymous):

is it 81?

OpenStudy (anonymous):

How did you get that?

OpenStudy (shamil98):

take the coefficient of b and divide it by 2 and then square it to get your value for c.

OpenStudy (shamil98):

18/2 = 9 9^2 = 81 x^2 + 18 + 81 = 0 is (x+9)^2

OpenStudy (anonymous):

Thank you so much that helped a lot!

OpenStudy (anonymous):

awesome i wasnt sure if i did it right. thanks shamil